OpenStudy (anonymous):

Find the slope of the line passing through the points (0, –5) and (4, 2)

OpenStudy (anonymous):

7/4

OpenStudy (igreen):

@tianalovesyou17 Don't give out answers.

OpenStudy (igreen):

To find the slope plug the two points into the equation: \(m = \dfrac{y_2-y_1}{x_2-x_1}\). Plug them in: \(m = \dfrac{2+5}{4-0}\) Subtract: \(m = \dfrac{5 + 2}{4}\) What's 5 + 2? @MikeBoi